# A Study on Adversarial Robustness of Discriminative Prototypical Learning

### v1

@@ Abstract
Deep neural networks demonstrate significant vulnerability to adversarial perturbations, posing risks for critical applications. Current adversarial training methods predominantly focus on robustness against attacks without explicitly leveraging geometric structures in the latent space, usually resulting in reduced accuracy on the original clean data. To address these issues, we propose a novel adversarial training framework named Adversarial Deep Positive-Negative Prototypes (Adv-DPNP), which integrates disriminative prototype-based learning with adversarial training. Adv-DPNP uses unified class prototypes serving dual roles as classifier weights and robust anchors, enhancing both intra-class compactness and inter-class separation in the latent space. Moreover, a novel dual-branch training mechanism maintains stable prototypes by updating them exclusively with clean data; while the feature extractor layers are learned using both clean and adversarial data to remain invariant against adversarial perturbations. In addition, our approach utilizes a composite loss function combining positive prototype alignment, negative prototype repulsion, and consistency regularization to further enhance discrimination, adversarial robustness, and clean accuracy. Extensive experiments conducted on standard benchmark datasets confirm the effectiveness of Adv-DPNP compared to state-of-the-art methods, achieving higher clean accuracy and competitive robustness under adversarial perturbations and common corruptions. Our code is available at https://github.com/fum-rpl/adv-dpnp
@@
adversarial training, adversarial robustness, discriminative prototypical learning, positive-negative prototypes.
@@ 1 Introduction
Deep neural networks (DNNs) have achieved remarkable performance in a wide range of tasks, including image classification [1, 2], speech recognition [3], and natural language processing [4]. However, these successes come with a critical vulnerability to adversarial attacks. As demonstrated in [5], DNNs are highly susceptible to subtle and carefully crafted perturbations added to their input, which can lead to a significant drop in accuracy. This vulnerability threatens the reliability and security of state-of-the-art neural models to be deployed in safety-critical applications such as autonomous driving [6], healthcare diagnostics [7], and security-sensitive systems [8], where ensuring robustness alongside high accuracy is essential.
@@
To mitigate the risks posed by adversarial vulnerabilities, various defense strategies have been developed by training models on adversarially perturbed examples [9, 10, 11, 12]. Despite their effectiveness, these approaches primarily focus on improving robustness against adversarial perturbations without explicitly considering the underlying geometric structure of the feature space. Therefore, they might suffer from reduced accuracy on original data and suboptimal generalization due to inadequate intra-class compactness and inter-class separation.
One promising direction to form a well-organized feature space is to utilize discriminative prototype-based learning [13]. Prototype-based methods represent each class with one or more prototypes in the feature space and classify samples based on their similarity to these learned prototypes. Such approaches not only offer an interpretable feature representation of classes but also inherently promote intra-class compactness and inter-class separation that are crucial for accurate classification. In our recent work [14], we introduced the Deep Positive-Negative Prototypes (DPNP) framework, which unifies the roles of classifier weights and class centers through shared prototype representations, thereby establishing a more structured latent space.
@@
Motivated by these observations, we propose a novel adversarial training framework named Adversarial Deep Positive-Negative Prototypes (Adv-DPNP), which extends our discriminative prototype-based learning paradigm to the adversarial domain. Adv‑DPNP is designed to enforce adversarial robustness into the learning process while preserving the concept of class prototypes and maintaining a well-structured latent space. In our approach, each class is represented by a positive prototype that serves as a robust anchor. These prototypes not only act as classifier weights but also shape compact and well-separated representations in the feature space. Building upon the unified design of DPNP, our method introduces a composite loss function that simultaneously aligns features with their corresponding positive prototype and enforces each positive prototype to remain sufficiently distant from the prototype of the nearest rival class, treated as a negative prototype. By pushing each positive prototype away from their counterpart negative prototypes, a sufficient margin between class prototypes is preserved.
@@
Additionally, we propose a novel dual-branch training mechanism that leverages both clean and adversarial data through two branches: a clean-data branch, which processes original samples and naturally updates both class prototypes and the feature extractor parameters, and an adversarial-data branch, which handles adversarially perturbed data and only updates the feature extractor parameters while keeping prototypes unchanged. This dual-branch approach ensures that the learned prototypes capture the distribution of clean data, while the feature extractor learns robust representations aligned with both clean and adversarial inputs.
@@
  • We propose Adv-DPNP, a novel framework that integrates discriminative prototype-based learning with adversarial training. Adv-DPNP leverages class prototypes explicitly as classifier weights and robust anchors, resulting in improved intra-class compactness and inter-class separation in adversarial settings.
@@
Through extensive experiments conducted on widely used benchmarks, CIFAR-10 [15] and CIFAR-100 [15], we demonstrate that Adv‑DPNP consistently outperforms state-of-the-art adversarial training methods in terms of clean accuracy and robustness. Moreover, we provide an in-depth analysis of the discriminative quality of learned feature representations using several quantitative metrics, highlighting the effectiveness of Adv-DPNP in maintaining compactness and clear separation in latent space. The rest of this paper is organized as follows: Section 2 reviews related work on adversarial attacks, adversarial training defenses, and discriminative prototype-based learning. In Section 3, we present the proposed Adv-DPNP framework, detailing our dual-branch training strategy, unified prototype design, and composite loss functions. Section 4 provides our experimental results, comparisons with state-of-the-art methods, and detailed analysis of feature-space discriminability. Finally, Section 5 concludes the paper and discusses directions for future research.
@@ 2 Related Work
The field of adversarial robustness has been extensively studied, leading to the development of numerous attack and defense mechanisms [16]. This section provides an overview of related works, categorizing them into three key areas. First, we review the evolution of adversarial attacks, from early discoveries to state-of-the-art methods designed to challenge DNNs. Next, we discuss major defense strategies, including adversarial training and various regularization techniques. Finally, we explore discriminative prototype-based approaches that aim to enhance both interpretability and robustness by leveraging structured feature representations. These insights collectively establish the foundation for our proposed method, which integrates adversarial training with discriminative prototype-based learning to improve clean accuracy while maintaining robustness against adversarial perturbations and common corruptions.
@@ 2.1 Adversarial Attacks
Other notable methods include DeepFool [19], which seeks the minimal perturbation needed to cross the decision boundaries, and the Carlini and Wagner (C&W) attack [20], which formulates adversarial perturbation generation as an optimization problem with explicit distance constraints. More recently, Croce and Hein [21] introduced AutoAttack (AA), an ensemble of parameter-free attacks that combines multiple strategies, including PGD variants and other specialized attacks, into a single automated framework. Its design ensures that no single type of attack is overlooked, making it a robust and standardized tool for benchmarking adversarial defenses.
@@ 2.2 Defense Strategies
Following the significant advances in adversarial attack methods, another line of research has focused on defense strategies. One of the most influential defense approaches is Adversarial Training (AT) [9, 10], in which models are explicitly trained on adversarially perturbed inputs. Such methods adopt a min-max optimization procedure, where the inner maximization generates worst-case perturbations that maximize the model’s loss, while the outer minimization updates the model parameters to be resilient against these perturbations.
Subsequent methods such as TRADES [11] extend AT by introducing a regularization term that minimizes the divergence between the predictions on clean and adversarial examples, thus encouraging the model to balance the trade-off between clean accuracy and adversarial robustness. Furthermore, AT-HE [22] integrates AT with hypersphere embedding by normalizing features in the penultimate layer and weights in the softmax layer with an additive angular margin. Additional methods, such as Adversarial Logit Pairing [23] and feature scattering techniques [24], aim to enforce similarity between feature representations of clean and adversarial samples. Moreover, some studies have suggested to treat misclassified adversarial samples differently. Specifically, MART [12] emphasizes misclassified adversarial samples by minimizing the distributional difference between these adversarial samples and their clean counterparts. Similarly, A3T [25] generates adversarial samples differently depending on whether a sample is correctly classified or misclassified.
@@ 2.3 Prototype-Based learning and Discriminative Approaches
More recently, prototype-based representations have been integrated with DNNs to enhance discriminative power, interpretability, and generalization. For example, Prototypical Networks [29] compute class prototypes as the mean of feature representations of samples belonging to the same class within the latent space. Other related methods using discriminative loss functions, such as Center Loss (CL) [30] and Constrained Center Loss (CCL) [31], explicitly minimize the distances between feature representations and their corresponding class centers. Although these methods enhance the intra-class compactness of learned feature representations, they maintain separate parameter sets and rely on independent update mechanisms for classifier weights and prototypes, resulting in misalignment between these components and consequently insufficient inter-class separation. In response to these limitations, we recently proposed the DPNP framework [14], a unified representation that merges classifier weights and class prototypes into shared parameters. By unifying classifier weights with class prototypes, decision boundaries naturally align with the learned prototypes, enhancing intra-class compactness and inter-class separability and leading to a well-structured feature space. Furthermore, this approach eliminates the need to store separate vectors for the class prototypes and classifier weights, thereby preventing potential mismatches between these components [32].
@@
Despite the success of both adversarial training and discriminative prototype-based learning independently, the integration of these two paradigms remains relatively underexplored. Although standard adversarial training methods effectively improve robustness, they typically neglect the geometric structure of the latent space. As a consequence, the learned features may not be optimally structured for clear class separation, which can limit interpretability and even potentially reduce performance on clean data. To address these issues, several studies have attempted to investigate the role of discriminative loss functions such as CL and triplet loss in improving adversarial robustness [33, 34]. Similarly, methods proposed in [35, 36] have further expanded this concept by applying geometric constraints, including hyperspherical or convex polytope constraints, on intermediate network layers to enhance intra-class compactness and inter-class separation. However, these methods incorporate adversarial examples directly into the prototype update process and rely on separate parameter sets for classifier weights and class prototypes. Consequently, these design choices can negatively impact prototype stability and ultimately limit their ability to achieve robustness comparable to state-of-the-art defense, particularly under white-box attack scenarios where an adversary has complete access to model gradients.
Motivated by these findings, we propose Adv-DPNP, an extension of the DPNP framework explicitly designed to integrate discriminative prototype-based learning into the adversarial domain. In the next section, we introduce the Adv-DPNP framework in detail, describing our shared prototype representation, the composite loss function, and the dual-branch training mechanism.
@@ 3.1 Notations and Problem Formulation
In conventional DNN classifiers, the output logits for each class $j$ are computed as the inner product $w_{j}^{T}f(x;\theta)$, where $w_{j}\in\mathbb{R}^{d}$ is the weight vector associated with class $j$. Based on the unified prototype-based design of our recent work [14], we associate each class $j$ with a learnable parameter vector $c_{j}\in\mathbb{R}^{d}$, which simultaneously serves as both the class prototype and the classifier weight. This shared parameterization explicitly aligns the decision boundaries defined by the classifier weights with the geometric structure of class representations in the feature space. Consequently, these unified class prototypes act as shared anchors, serving both as classifier weights and as geometric centers for their respective classes in the latent space. To effectively shape the geometry of the latent space, our prototypes $c_{j}$ are constrained to lie on the surface of a hypersphere with radius $\alpha$. This constraint allows them to actively influence the organization of feature representations rather than being passively estimated as extra separate parameters, usually through mean calculations [31].
In our proposed framework, the probability $p_{j}(x)$ assigned to class $j$ for input $x$ is computed using a softmax function over the similarity scores between the feature representation $f(x;\theta)$ and the prototype $c_{j}$ relative to other class prototypes $c_{k}$:
@@
where $p_{y_{i}}(x_{i})$ denotes the predicted probability assigned to the true class $y_{i}$ for the sample $x_{i}$.
3.2 Composite Loss Function
DPP Loss. Building upon the unified prototypes ${c_{j}}$, we define the DPP loss that integrates two components, the CE loss for accurate classification and a positive prototype alignment term, encouraging the feature representations $f(x_{i};\theta)$ to closely align with their corresponding class prototypes $c_{y_{i}}$:
@@ 3.2 Composite Loss Function
where $c_{y_{i}}$ is the true class prototype and $\lambda_{\text{DPP}}$ controls the trade-off between correct classification and clustering around each class prototype $c_{y_{i}}$. By minimizing the Euclidean distance between $f(x_{i};\theta)$ and the corresponding positive prototype $c_{y_{i}}$, the features are pulled closer to the correct class center, promoting intra-class compactness in the latent space.
@@
This term is sample-independent and is computed separately for each class. Hence, it is applied once per mini-batch, according to the positive prototype update rate. The chosen $L_{1/2}$ norm exerts stronger repulsion when the centers are relatively close, effectively increasing the separations among the nearest classes. As a result, DNP promotes inter-class separability in the latent space.
DFA Loss. While DPP and DNP collectively improve classification and impose geometric constraints on prototypes, they do not explicitly ensure consistency between clean and adversarial representations derived from feature extractor output. We incorporate the DFA loss $L_{\text{DFA}}(x_{i},\tilde{x_{i}})$ derived from the Kullback-Leibler divergence, similar to TRADES [11], which penalizes discrepancies in predicted distributions for clean inputs $x_{i}$ and their adversarial counterparts $\tilde{x_{i}}$:
@@
where $\lambda_{\text{DNP}}$ and $\lambda_{\text{DFA}}$ are hyperparameters determining the relative contributions of DNP and DFA losses. This composite formulation balances intra-class compactness (DPP), inter-class separability (DNP), and distribution consistency (DFA), collectively enhancing both clean data accuracy and adversarial robustness.
Although adversarial examples $\tilde{x_{i}}$ also pass through the same set of loss terms as original inputs $x_{i}$, in our framework the prototypes $c_{j}$ only receive gradient updates from clean data. This design choice stabilizes the prototype geometry by preventing adversarial perturbations from shifting class centers. In the next section (Section 3.3), we elaborate on this dual-branch training procedure and explain how updates are decoupled for clean and adversarial data.

### v2

# Deep Positive-Negative Prototypes for Adversarially Robust Discriminative Prototypical Learning

@@ Abstract
Despite the advantages of discriminative prototype-based methods, their role in adversarial robustness remains underexplored. Meanwhile, current adversarial training methods predominantly focus on robustness against adversarial attacks without explicitly leveraging geometric structures in the latent space, usually resulting in reduced accuracy on the original clean data. We propose a novel framework named Adversarially trained Deep Positive-Negative Prototypes (Adv-DPNP), which integrates discriminative prototype-based learning with adversarial training. Adv-DPNP uses unified class prototypes that serve as both classifier weights and robust anchors in the latent space. Moreover, a novel dual-branch training mechanism maintains stable prototypes by updating them exclusively with clean data, while the feature extractor is trained on both clean and adversarial inputs to increase invariance to adversarial perturbations. In addition, we use a composite loss that combines positive-prototype alignment, negative-prototype repulsion, and consistency regularization to further enhance discrimination, adversarial robustness, and clean accuracy. Extensive experiments on standard benchmarks (CIFAR-10/100 and SVHN) confirm that Adv-DPNP improves clean accuracy over state-of-the-art defenses and baseline methods, while maintaining competitive or superior robustness under a suite of widely used attacks, including FGSM, PGD, C&W, and AutoAttack. We also evaluate robustness to common corruptions on CIFAR-10-C, where Adv-DPNP achieves the highest average accuracy across severities and corruption types. Additionally, we provide an in-depth analysis of the discriminative quality of the learned feature representations, highlighting the effectiveness of Adv-DPNP in maintaining compactness and clear separation in the latent space.
@@
discriminative prototypical learning , positive-negative prototypes , adversarial robustness , adversarial training
[…]
[1]organization=Computer Engineering Department, Ferdowsi University of Mashhad, city=Mashhad, country=Iran
@@ 1 Introduction
Prototype-based learning has long been recognized as an interpretable and geometrically principled approach for multiclass classification by representing each class with one or more prototypes [50, 40, 22]. However, these methods historically have been less competitive in natural (clean) accuracy compared to modern deep neural networks (DNNs), which dominate tasks such as image classification [9, 15], speech recognition [14], and natural language processing [7]. Recently, discriminative extension of prototype-based learning is proposed which closes the accuracy gap with DNNs, while offering compact and well-separated decision regions [49]. However, DNNs remain highly susceptible to carefully crafted adversarial perturbations added to their input, which can lead to a significant drop in accuracy [41]. This vulnerability is particularly critical in safety-sensitive applications such as autonomous driving [18], medical diagnosis [11], and security-sensitive systems [27], where models must sustain both high accuracy and robustness.
@@
To mitigate the risks posed by adversarial vulnerabilities, various defense strategies have been developed by training models on adversarially perturbed examples [13, 30, 52, 42, 6]. Despite their effectiveness, these approaches primarily focus on improving robustness against adversarial perturbations without explicitly considering the underlying geometric structure of the feature space. Therefore, they might suffer from reduced clean accuracy on the original data and suboptimal generalization due to inadequate intra-class compactness and inter-class separation. In such Adversarial Training (AT) approaches, the gradients from perturbed samples may detrimentally shift decision boundaries, leading to reduced stability and discrimination.
Accordingly, some recent works have explored using prototypes and geometric constraints to address these issues [46, 10, 34, 33]. Nevertheless, these approaches either fix classifier weights, which limits the model’s ability to adapt margins to data complexity, or keep classifier weights and prototypes as separate parameter sets, leading to misalignment and unstable decision boundaries [20]. Another issue is updating prototypes using adversarial samples, which contaminates the prototypes with adversarial noise and reduces their stability.
@@
Motivated by these observations, we propose a novel framework named Adversarially trained Deep Positive-Negative Prototypes (Adv-DPNP), which extends our discriminative prototype-based learning paradigm [49] to the adversarial domain. Adv‑DPNP is designed to enforce adversarial robustness into the learning process while preserving the concept of class prototypes and maintaining a well-structured latent space. In our approach, each class is represented by a positive prototype that serves as a robust anchor and acts as classifier weights for the corresponding class at the same time, while the nearest rival class centers are treated as negative prototypes. Hence, a compact and well-separated feature space is formed, which combines the benefits of prototype-based methods and discriminative learning.
@@
Additionally, we propose a novel dual-branch training mechanism that leverages both clean and adversarial data through two separate branches: the clean-data branch processes original samples and updates both class prototypes and the feature extractor parameters, while the adversarial-data branch handles adversarially perturbed data and only updates the feature extractor parameters and keeps prototypes unchanged. This dual-branch approach ensures that the learned prototypes capture the distribution of clean data, while the feature extractor learns robust representations by encouraging the alignment of corresponding clean and adversarial inputs in the latent space.
@@
  • 1.
    We propose Adv-DPNP, a novel framework that integrates discriminative prototype-based learning with adversarial training. Adv-DPNP explicitly unifies classifier weights and prototypes into shared parameters, ensuring that decision boundaries and geometric class centers remain inherently aligned in adversarial settings.
  • 2.
@@
Through extensive experiments conducted on widely used benchmarks, CIFAR-10/100 [24] and SVHN [35], we demonstrate that Adv‑DPNP consistently outperforms state-of-the-art (SOTA) adversarial training methods in terms of clean accuracy and robustness. Also, we provide an in-depth analysis of the discriminative quality of learned feature representations using several quantitative metrics, highlighting the effectiveness of Adv-DPNP in maintaining compactness and clear separation in latent space. The rest of this paper is organized as follows: Section 2 reviews related work on adversarial attacks, adversarial training defenses, and discriminative prototype-based learning. In Section 3, we present the Adv-DPNP framework, detailing our dual-branch training strategy, unified prototype design, and composite loss functions. Section 4 provides experimental results, comparisons with SOTA methods, and a detailed analysis of feature space discriminability. Finally, Section 5 concludes the paper and discusses directions for future research.
@@ 2 Related Work
The field of adversarial robustness has been extensively studied, leading to the development of numerous attack and defense mechanisms [43]. This section provides an overview of related works, categorizing them into three key areas. First, we review the evolution of adversarial attacks, designed to challenge DNNs. Next, we discuss major defense strategies, including adversarial training and various regularization and feature-manipulation techniques. Finally, we explore discriminative prototype-based approaches that aim to enhance both interpretability and robustness by leveraging structured feature representations. These insights collectively establish the foundation for our proposed method, which integrates adversarial training with discriminative prototype-based learning to improve clean accuracy while maintaining robustness against adversarial perturbations and common corruptions.
@@ 2.1 Adversarial Attacks
Other notable methods include DeepFool [32], which seeks the minimal perturbation needed to cross the decision boundaries, and the Carlini and Wagner (C&W) attack [3], which formulates adversarial perturbation generation as an optimization problem with explicit distance constraints. More recently, Croce and Hein [5] introduced AutoAttack (AA), an ensemble of parameter-free attacks including PGD variants, a targeted decision-boundary attack, and a black-box attack. Its design ensures that no single type of attack is overlooked, making it a robust and standardized tool for benchmarking adversarial defenses.
@@ 2.2 Defense Strategies
Following the significant advances in adversarial attack methods, another line of research has focused on defense strategies. One of the most influential defense approaches is AT [13, 30], in which models are explicitly trained on perturbed inputs. Such methods adopt a min-max optimization procedure, where the inner maximization generates worst-case perturbations that maximize the model’s loss, while the outer minimization updates the model parameters to be resilient against these perturbations.
Subsequent methods, such as AT-HE [36], integrate AT with hypersphere embedding by normalizing features in the penultimate layer and weights in the softmax layer with an additive angular margin. Additional methods, such as Adversarial Logit Pairing [19] and feature scattering techniques [51], aim to enforce similarity between feature representations of clean and adversarial samples. Moreover, MART [42] specifically treats misclassified adversarial samples by minimizing the distributional discrepancy between them and their clean counterparts, thereby focusing training on hard examples. Furthermore, TRADES [52] extends AT by adding a Kullback–Leibler (KL) regularization term that minimizes the divergence between the predictions on clean and adversarial examples, thus encouraging the model to balance the trade-off between clean accuracy and adversarial robustness. Building on the same idea, the Improved KL (IKL) loss [6] reformulated the asymmetric KL objective as the sum of a weighted mean-squared error and a soft-label cross-entropy; hence eliminating the optimization asymmetry leading to better class-wise global statistics and mitigation of sample-wise bias.
A separate line of work argues that the weak accuracy–robustness trade-off arises because vanilla AT excessively increases the decision-boundary margin. Helper-based Adversarial Training (HAT) [38] addresses this by adding helper samples (extrapolated adversarial images given soft labels from a standard model) to pull the boundary back toward informative directions, boosting clean accuracy with negligible robustness loss. Follow-up schemes pursue the same intuition, Increasing-Margin Adversarial Training [29] adaptively enlarges or shrinks margins through an equilibrium analysis and has been adopted in medical-imaging pipelines. Analytic tools such as ATVis [53] later visualized these margin dynamics and confirmed HAT’s findings. Our Adv-DPNP regularizes the margins via prototypes and discriminative terms in its loss function and therefore remains compatible with these works.
Beyond loss-level regularization, several researchers attempt to suppress or adapt internal feature activations corrupted by adversarial noise. Feature Denoising [45] is proposed to denoise intermediate maps via non-local means and bilateral filtering, while Channel Activation Suppression [2] zeros out non-robust channels during training. CIFS [47] goes further by selecting channels according to a learned importance mask that suppresses less-relevant features. Most recently, Kim et al. introduced Feature Separation and Recalibration (FSR) [21], which first separates robust and non-robust activations via a Gumbel-Softmax learnable mask and then recalibrates the non-robust branch instead of discarding it. In contrast to these activation-level methods, our Adv-DPNP keeps all activations but organizes the latent space through prototype anchors; thereby, it can be complemented with such feature-manipulation modules without extra architectural overhead.
@@ 2.3 Prototype-Based learning and Discriminative Approaches
More recently, prototype-based representations have been integrated with DNNs to enhance discriminative power, interpretability, and generalization. For example, Prototypical Networks [40] compute class prototypes as the mean of feature representations of samples belonging to the same class within the latent space. Other related methods using discriminative loss functions, such as Center Loss (CL) [44, 39], explicitly minimize the distances between feature representations and their corresponding class centers. Although these methods enhance the intra-class compactness of learned feature representations, they maintain separate parameter sets and rely on independent update mechanisms for classifier weights and prototypes, resulting in misalignment between these components and consequently insufficient inter-class separation [49, 20]. In response to these limitations, we recently proposed the DPNP framework [49], a unified representation that merges classifier weights and class prototypes into shared parameters. By unifying classifier weights with class prototypes, decision boundaries naturally align with the learned prototypes, enhancing intra-class compactness and inter-class separability and leading to a well-structured feature space. Furthermore, this approach eliminates the need to store separate vectors for the class prototypes and classifier weights, thereby preventing potential mismatches between these components [20].
@@
Despite the success of both discriminative prototype-based learning and adversarial training independently, the integration of these two paradigms remains relatively underexplored. Although standard adversarial training methods effectively improve robustness, they typically neglect the geometric structure of the latent space. As a consequence, the learned features may not be optimally structured for clear class separation, which can limit interpretability and potentially reduce performance on clean data. To address these issues, several studies have attempted to investigate the role of discriminative loss functions such as CL and triplet loss in improving adversarial robustness [31, 17]. Similarly, methods proposed in [34, 33] have further expanded this concept by applying geometric constraints, including hyperspherical or convex polytope constraints, on intermediate network layers to enhance intra-class compactness and inter-class separation. However, these methods incorporate adversarial examples directly into the prototype update process and rely on separate parameter sets for network weights and prototypes, which can lead to misalignment between them. Consequently, these design choices can negatively impact prototype stability [20] and ultimately limit their ability to achieve robustness comparable to SOTA defense, particularly under white-box attack scenarios where an adversary has complete access to model gradients.
An alternative line of work tackles this misalignment by fixing or regularizing the classifier weights themselves. Orthogonal Classifier (OC) [46] sets the softmax layer to a precomputed, fully dense orthogonal matrix with equal magnitude entries, enforcing uniform angular separations for robustness improvement. Building on this idea, Diversity via Orthogonality (DIO) [10] augments the network with multiple parallel classifier heads, then injects layer-wise orthogonality regularizers together with a margin-maximization loss to promote feature diversity and robustness. These orthogonality-based schemes enlarge decision margins through fixed geometric constraints rather than learnable prototypes. Unlike OC and DIO, which fix or strictly regularize the classifier weights, our Adv-DPNP keeps the prototypes trainable, allowing flexible margin adjustment while preserving classifier weights-prototypes alignment.
@@ 3.1 Notations and Problem Formulation
In conventional DNN classifiers, the output logits for each class $j$ are computed as the inner product $w_{j}^{T}f(x;\theta)$, where $w_{j}\in\mathbb{R}^{d}$ is the weight vector associated with class $j$. Based on the unified prototype-based design of our recent work [49], we associate each class $j$ with a learnable parameter vector $c_{j}\in\mathbb{R}^{d}$, which simultaneously serves as both the class prototype and the classifier weight. This shared parameterization explicitly aligns the decision boundaries defined by the classifier weights with the geometric structure of class representations in the feature space. Consequently, these unified class prototypes act as shared anchors, serving both as classifier weights and as geometric centers for their respective classes in the latent space.
3.2 Composite Loss Function
DPP Loss. To effectively shape the geometry of the latent space, our prototypes $c_{j}$ are constrained to lie on the surface of a hypersphere with radius $\alpha$, playing a dual role as class centers (point representatives in the latent space) and as the weight vector of neuron $j$ in the classifier. This constraint allows them to actively influence the organization of feature representations rather than being passively estimated as extra separate parameters, usually through mean calculations [49]. In our proposed framework, the probability $p_{j}(x)$ assigned to class $j$ for input $x$ is computed using a softmax function over the similarity scores between the feature representation $f(x;\theta)$ and the prototype $c_{j}$ relative to other class prototypes $c_{k}$:
@@ 3.2 Composite Loss Function
where $p_{y_{i}}(x_{i})$ denotes the predicted probability assigned to the true class $y_{i}$ for the sample $x_{i}$. Building upon the unified prototypes $c_{j}$}, we define the DPP loss that integrates two components, the CE loss for accurate classification and a positive prototype alignment term, encouraging the feature representations $f(x_{i};\theta)$ to closely align with their corresponding class prototypes $c_{y_{i}}$:
@@
where $c_{y_{i}}$ is the true class prototype and $\lambda_{\text{DPP}}$ controls the trade-off between correct classification and clustering around each class prototype $c_{y_{i}}$. By minimizing the Euclidean distance between $f(x_{i};\theta)$ and the corresponding positive prototype $c_{y_{i}}$, the features are pulled closer to the correct class center, promoting intra-class compactness in the latent space. In addition, at the start of each epoch, all prototypes $c_{j}$ are renormalized to keep them on a hypersphere of radius $\alpha$ (i.e., $\|c_{j}\|_{2}=\alpha$); hence, $1/\alpha$ acts as a temperature scaling factor that helps to obtain well-behaved CE gradients and improve numerical stability.
@@
This term is sample-independent and is computed individually for each class. Hence, it is applied once per mini-batch, according to the positive prototype update rate. We select the norm $\ell_{1/2}$ to emphasize short-range repulsion between nearest prototypes. This exerts stronger repulsion when the centers are relatively close, effectively increasing the separations among the nearest classes.
[…]
@@
where $\lambda_{\text{DNP}}$ and $\lambda_{\text{DFA}}$ are hyperparameters determining the relative contributions of DNP and DFA losses. This composite formulation implements a prototype-based learning framework which generalizes AT/TRADES losses. Here DPP serves as the classification objective and explicitly induces intra-class compactness; DNP acts as a margin regularizer at the prototype level, whereas DFA enforces clean–adversarial distributional consistency without architectural overhead. The result is a geometrically well-structured latent space.
It should be noted that in our framework, although adversarial examples $\tilde{x_{i}}$ pass through the same set of loss terms as clean inputs $x_{i}$, the prototypes $c_{j}$ only receive gradient updates from the clean data. This design choice stabilizes the prototype geometry by preventing adversarial perturbations from shifting class centers. In the next section (Section 3.3), we elaborate on this dual-branch training procedure and explain how updates are decoupled for clean and adversarial data.